An electron is moving with a kinetic energy of 0.923 MeV. What is its speed?

  1. An electron is moving with a kinetic energy of 0.923 MeV. What is its speed?
Answer #1

Q. Answer energy Since we kneen that the test elechon moc² = 0.51 Mev 0.923 Mevz Noud, KE = (x-1) mc² - 17-1) = KE - 0.923 Meplease like thank you....